# Multi-Level Feature Aggregation and Recursive Alignment Network for Real-Time Semantic Segmentation

### v1

@@ II-C Feature Alignment/Upsampling
DCNNs need to enlarge the receptive field through repeated max-pooling or stride convolution, so the high-level features of the encoder and the final predictions are inevitably downsampled. Bilinear interpolation is widely used to upsample feature/score maps [32, 11, 13], as it is simply and fast. However, it is oversimple and data-independent, making it difficult to align multi-level features. To avoid the spatial misalignment, Noh et al. [39] adopted unpooling layers and deconvolution layers to build the decoder to recover the details of low-resolution feature maps. To ensure efficiency, SegNet [40] uses the pooling positions stored in the encoder to achieve nonlinear upsampling of the corresponding level features of the decoder. Recently, GUN [29] proposes a transformation module to learn 2D offsets to guide upsampling between two different resolution features for real-time segmentation. AlignSeg [41], FaPN [42] and SFNet [36] share a similar motivation and all propose using flow-based alignment module to estimate 2D transformation offsets to implement learnable interpolation. They design the alignment module to fuse features between different levels of the encoder, so they need to bridge the semantic gap when performing spatial alignment. The semantic gap between multi-level features may mislead the transform offset learning in the alignment module.
III Our Proposed Real-time Approach
III-A Overview
@@
As shown in Fig. 1, three complementary modules are proposed to build our network: MFAM, RAM and ASFM. For MFAM, we build it through top-down and bottom-up paths, as well as the lateral connections for information exchange between paths. Such a structure has three advantages. Firstly, compared to long-range skip connections, it can aggregate long-distance features in an information flow manner to narrow the semantic gap. Secondly, by aggregating hierarchical features of the encoder to each scale, this architecture can provide both high-level semantic information and low-level spatial details for the prediction of the corresponding scale. Thirdly, since each feature map from MFAM aggregates multi-level features of the same encoder, we have reason to believe that the semantic gap between the output features is largely narrowed. This will benefit the transform offset learning in the following spatial alignment module. Upon the multi-scale features obtained from MFAM, we build four independent paths to perform multi-scale inference. Before ASFM, we first design RAM to combine flow-based alignment module with the recursive upsampling architecture to address the spatial misalignment between multi-scale score maps. In the end, we propose ASFM to do adaptive fusion over multi-scale scores, and the training is carried out by our customized supervision function MJS.
III-B Multi-level Feature Aggregation Module (MFAM)
Figure 2: The details of “Feature Fusion” module used in Bottom-up path or Top-down path. (a) In Bottom-up path. (b) In Top-down path.
Figure 3: The comparison between different multi-level feature aggregation methods. ”Aux Loss” denotes auxiliary loss. The dashed boxes indicate that the corresponding components are discarded during the inference phase. (a) FPN-like network. (b) LSCA (our initial design). (c) Our MFAM.
@@ III-B Multi-level Feature Aggregation Module (MFAM)
We start with the encoder to describe how to build our MFAM. There are three reasons for choosing ResNet-18 as the backbone to obtain multi-level features. First, it is a good fit for fine-tuning since the pre-trained parameters are publicly available. Our network can benefit from knowledge transfer by loading their weights trained on imageNet. Second, it promotes training efficiency due to moderate depth and residual structure. Third, it has low model complexity and can be used to construct real-time segmentation networks. The ResNet-18 has four stages (a.k.a., blocks), which extract multi-level feature maps with spatial resolutions of 1/4, 1/8, 1/16, and 1/32 of the input image (${S_{1}}\sim{S_{4}}$ in Fig. 1(a)). Additionally, we use two atrous convolutional layers [32] with dilation rate of 2 and 4 to replace the stride convolutions in the last two stages of the backbone to enlarge receptive field. Differently, we do not change the stride length of stride convolutions from 2 to 1 as done in [32, 11], thus increasing the receptive field of the last two stages without changing the feature size (still 1/16 and 1/32 of the input image). This can significantly increase accuracy with a slight decrease in inference speed (Table. I, native FPN vs. improved FPN).
@@
Upon multi-level features of the encoder, we flow the topmost-level feature to the lowest-level feature through the top-down path. In this process, we utilize feature fusion modules to aggregate mid-level features and pass them to the lower-level features, so that the outputs of top-down path fuse higher-level features (e.g., ${U_{2}}$ aggregates ${S_{2}}\sim{S_{4}}$). It is worth noting that all the features from the encoder use a 1$\times$1 convolutional layer to adjust the channel dimension before inputting the top-down path or bottom-up path, which is omitted in Fig. 1(a) for brevity. For the feature fusion modules in top-down path (Fig. 2), in order to ensure efficiency, we use bilinear interpolation to upsample higher-level features, and then use element-wise addition followed by a 3$\times$3 convolutional layer to fuse features. This process can be expressed mathematically as:
@@
Here, we directly get ${U_{4}}$ from ${S_{4}}$: ${U_{4}}={\rm{con}}{{\rm{v}}_{1\times 1}}({S_{4}})$. ${\rm{con}}{{\rm{v}}_{1\times 1}}$ represents a 1$\times$1 convolutional layer for feature channel adjustment. ${\rm{con}}{{\rm{v}}_{3\times 3}}$ is a 3$\times$3 convolutional layer for feature fusion. $\rm{Up}$ means bilinear upsampling. Similarly, we construct a bottom-up path to flow the lowest-level feature in the encoder to the topmost feature, so that the resulting features aggregate the lower-level features. We still employ a lightweight feature fusion module to fuse adjacent-level features (Fig. 2), but we use bilinear interpolation to downsample features instead of upsampling. This can be mathematically expressed as:
@@
We show a comparison of our MFAM with FPN-like architecture [24, 18, 36] and Long-range Skip Connection based Architecture (LSCA) in Fig. 3 to give readers a clearer understanding of the motivations of our design. FPN-like architecture utilizes the top-down path with lateral connections to obtain high-level semantic features at all stages, while its topmost and mid-scale features lack low-level information. The LSCA is the one we originally designed for fusing multi-level features. Although it could fuse multi-level features in an explicit manner, using the long skip connections between long-distance features (e.g., ${S_{4}}$ to ${S_{1}}$) for fusion will lead to a larger semantic gap and misalignment than fusion between adjacent features. Quantitative comparisons are given in Table. II.
III-C Recursive Alignment Module (RAM)
Inspired by [41, 42, 36], we adopt a flow-based alignment module in RAM to learn 2D transformation offsets to guide the upsampling between features with different resolutions. However, we have lots of noticeable difference from above methods. Firstly, the above methods use this flow-based alignment method to bridge the semantic gap to fuse multi-level features of the encoder, while our RAM aims to achieve spatial alignment between score maps by aligning features before the segmentation head. Secondly, the previous MFAM can largely narrow the semantic gap betweent multi-scale feautres, so our module can perform pure spatial alignment without being disturbed by the semantic gap. Thirdly, we align features in a stepwise manner.
@@ III-C Recursive Alignment Module (RAM)
In Fig. 4, we take the alignment between ${F_{4}}$ and ${F_{1}}$ as an example to compare our proposed RAM with the straightforward alignment method. As shown in Fig. 4, the straightforward method puts ${F_{4}}$ and ${F_{1}}$ into the flow-based alignment module to learn the corresponding offset, and then use the alignment function to directly align ${F_{4}}$ to ${F_{1}}$ without using intermediate features. Compared with the straightforward alignment, our designed RAM is not only more accurate but also more efficient. Here, we give computational complexity analysis. Since the complexity of other operations in RAM and the straightforward alignment method is negligible compared to the 3 $\times$ 3 convolutional operation, the overall complexity can be estimated by 3 $\times$ 3 convolutional operations. Therefore, the complexity of computing the offset between ${F_{4}}$ and ${F_{3}}$ can be expressed as $W\times H\times{C_{in}}\times{C_{out}}\times 3\times 3$, which can be simplified as $9WH{C^{2}}$. Here, $H$ and $W$ denote the height and width of ${F_{3}}$, and ${C_{in}}$ and ${C_{out}}$ represent the channels of input and output features of the convolutional layer. Since the height and width of ${F_{1}}$ are increased by 4 times compared to ${F_{3}}$, the complexity of inputting upsampled ${F_{4}}$ and ${F_{1}}$ into the 3 $\times$ 3 convolutional layer can be expressed as $4W\times 4H\times{C_{in}}\times{C_{out}}\times 3\times 3$. Ultimately, the ratio between the overall complexity of RAM and the straightforward alignment method is: $(9WH{C^{2}}+4\times 9WH{C^{2}}+16\times 9WH{C^{2}})/(3\times 16\times 9WH{C^{2}})=7/16$. In terms of accuracy, the quantitative comparison and corresponding analyses are given in Section. IV-B3.
III-D Adaptive Scores Fusion Module
@@
III-E Multi-scale Joint Supervision
@@ IV-A Training Protocol
TABLE I: Comparison with the baseline method. MSL: Multi-Stage Loss. MJS: Multi-scale Joint Supervision. OHEM: Online Hard Example Mining.
@@
TABLE II: Comparison with different multi-level feature aggregation methods. FPN-like: the FPN-like Network shown in Fig. 3. LSCA: the Long-range Skip Connection based Architecture in Fig. 3.
@@
TABLE III: Ablation study on RAM. Bilinear: bilinear interpolation. SA: Straightforward Alignment.
@@
TABLE IV: Comparison between three multi-scale scores fusion methods: Average, Maximum, and ASFM.
@@
TABLE V: Ablation study on importance of each component. ”-”: Indicates that the corresponding experiment directly utilizes the features of the encoder for alignment and prediction. Bilinear: bilinear interpolation. Cancat: the channel concatenation operation for feature fusion.
@@
TABLE VI: Ablation Study for booster training strategy. OHEM: Online Hard Example Mining.
@@
TABLE VII: Ablation for Multi-scale predictions. ”-”: The corresponding score map and weight map are discarded during training and inference phase. The proposed MJS + OHEM is used to supervise the training of each experiment.
@@
TABLE VIII: Experiments for the scale selection during the inference to further reduce model complexity. All ablations are trained on the complete architecture with four scales and discarded certain scales only during the inference phase. The results reaching the best trade-off are bolded.
@@
IV-B Ablation Studies
IV-B1 Comparison with the Baseline Methods
As shown in Fig. 3, we choose the FPN-like network as the baseline method and employs ResNet-18 as its backbone. For a fair comparison, we also adopt two atrous convolutional layers to the backbone of the baseline, following the way described in Section. III-B. Furthermore, we add three additional segmentation heads at different stages of the baseline to compute auxiliary loss to boost network training, which is named Multi-Stage Loss (MSL). Notably, by inferring on multi-scale features, our MFARANet can calculate MSL without adding additional segmentation head. As shown in Table. I, through the above operations, we improve the accuracy of the native FPN from 72.3% to 74.7% of the improved FPN with a slight decrease in inference speed. Using the same backbone and MSL, our method can achieve 77.0% mIoU, which is 2.3% higher than the strong baseline. At the same time, the computational complexity is reduced by 10.5 GFLOPs, achieving a faster inference speed. Furthermore, by combining MJS and OHEM to boost the training process, our method further improves the accuracy to 78.2%, which is 5.9% and 3.5% higher than native FPN and improved FPN, respectively. Qualitative segmentation comparison results between our method and the baseline on Cityscapes validation set are shown in Fig. 6. Our method shows improvement on multi-scale objects.
IV-B2 Ablation Study for MFAM
To demonstrate the superiority of MFAM, we compare it with the other two feature aggregation methods shown in Fig. 3: FPN-like architecture and LSCA. In Table. II, we replace the MFAM in our network with FPN and LSCA, respectively, and all other components remain unchanged. From the results, our method improves the accuracy by 0.7% and 1.6% compared to FPN and LSCA, respectively, with less than 6 GFLOPs increase in computational complexity. We attribute the superiority of our MFAM as two aspects: a) The topmost and mid-scale features of FPN lack low-level features, while each feature from MFAM aggregates full-level features of the encoder. b) We aggregate long-distance features in an information flow manner, which narrows the semantic gap compared to using long-range skip connections.
IV-B3 Ablation Study for RAM
We propose RAM to align scaled features and carry out several ablation studies to reveal its effectiveness. First, we utilize bilinear interpolation to replace RAM for upsampling, and the first row in Table. III has a decrease in accuracy of 1.5% compared to the third row. This shows that spatial misalignment is one of the key issues for multi-scale scores fusion and can be efficiently solved by our RAM. Second, we compare Straightforward Alignment (Fig. 4) with our proposed Recursive Alignment (Fig. 4), and our RAM improves the accuracy and inference speed by 0.5% and 2.6 FPS, respectively. We attribute this to adopting the stepwise alignment to replace direct alignment between long-distance features. Adjacent features have less spatial misalignment due to smaller scale differences compared to long-range features, leading to more accurate offsets learning. Computational complexity analysis has been given in Section. III-C.

### v3

@@ 2.3 Feature Alignment/Upsampling
DCNNs need to enlarge the receptive field through repeated max-pooling or stride convolution, so the high-level features of the encoder and the final predictions are inevitably downsampled. Bilinear interpolation is widely used to upsample feature/score maps [38, 10, 12], as it is simply and fast. However, it is oversimple and data-independent, making it difficult to align multi-level features. To avoid the spatial misalignment, Noh et al. [45] adopted unpooling layers and deconvolution layers to build the decoder to recover the details of low-resolution feature maps. To ensure efficiency, SegNet [46] uses the pooling positions stored in the encoder to achieve nonlinear upsampling of the corresponding level features of the decoder. Recently, GUN [35] proposes a transformation module to learn 2D offsets to guide upsampling between two different resolution features for real-time segmentation. AlignSeg [47], FaPN [48] and SFNet [42] share a similar motivation and all propose using flow-based alignment module to estimate 2D transformation offsets to implement learnable interpolation. The above flow-based alignment methods fuse features between different levels of the encoder, so they need to bridge the semantic gap when performing spatial alignment. The semantic gap between multi-level features may mislead the transform offset learning in the alignment module. Furthermore, they are used for adjacent features without large resolution differences, while our RAM is specifically designed for aligning long-distance feature maps.
3 Methods
3.1 Overview
@@
As shown in Fig. 3, three complementary modules are proposed to build our parallel inference network: MFAM, RAM and ASFM. For MFAM, we build it through top-down and bottom-up paths, as well as the lateral connections for information exchange between paths. Such a structure has three advantages. Firstly, compared to long-range skip connections (Fig. 5), it can aggregate long-distance features in an information flow manner to narrow the semantic gap. Secondly, by aggregating hierarchical features of the encoder to each scale, this architecture can provide both high-level semantic information and low-level spatial details for the prediction of the corresponding scale. Thirdly, since each feature map from MFAM aggregates multi-level features of the same encoder, we have reason to believe that the semantic gap between the output features is largely narrowed. This will benefit the transform offset learning in the following spatial alignment module. Upon the multi-scale features obtained from MFAM, we build four independent parallel paths to perform multi-scale inference. Before ASFM, we first design RAM to combine flow-based alignment module with the recursive upsampling architecture to address the spatial misalignment between multi-scale score maps. In the end, we propose ASFM to do adaptive fusion over multi-scale scores, and the training is carried out by our customized supervision function MJS.
3.2 Multi-level Feature Aggregation Module (MFAM)
@@
We start with the encoder to describe how to build our MFAM. There are two main reasons for choosing ResNet-18 as the backbone to obtain multi-level features. First, it is a good fit for fine-tuning since the pre-trained parameters are publicly available. Our network can benefit from knowledge transfer by loading their weights trained on imageNet. Second, due to the moderate depth and residual structure, it promotes training efficiency and has low model complexity, which can be used to build real-time segmentation networks. The ResNet-18 has four stages (a.k.a., blocks), which extract multi-level feature maps with spatial resolutions of 1/4, 1/8, 1/16, and 1/32 of the input image (${S_{1}}\sim{S_{4}}$ in Fig. 3(a)). Additionally, we use two atrous convolutional layers [38] with dilation rate of 2 and 4 to replace the stride convolutions in the last two stages of the backbone to enlarge receptive field. Differently, we do not change the stride length of stride convolutions from 2 to 1 as done in [38, 10], thus increasing the receptive field of the last two stages without changing the feature size (still 1/16 and 1/32 of the input image). This can significantly increase accuracy with a slight decrease in inference speed (Table. 1, native FPN vs. improved FPN).
@@
Upon multi-level features of the encoder, we flow the topmost-level feature to the lowest-level feature through the top-down path. In this process, we utilize feature fusion modules to aggregate mid-level features and pass them to the lower-level features, so that the outputs of top-down path fuse higher-level features (e.g., ${U_{2}}$ aggregates ${S_{2}}\sim{S_{4}}$). It is worth noting that all the features from the encoder use a 1$\times$1 convolutional layer to adjust the channel dimension before inputting the top-down path or bottom-up path, which is omitted in Fig. 3(a) for brevity. For the feature fusion modules in top-down path, we adopt a simple and general design to ensure efficiency. We use bilinear interpolation to upsample higher-level features, and then use element-wise addition followed by a 3$\times$3 convolutional layer to fuse features. This process can be expressed mathematically as:
@@
Here, we directly get ${U_{4}}$ from ${S_{4}}$: ${U_{4}}={\rm{con}}{{\rm{v}}_{1\times 1}}({S_{4}})$. ${\rm{con}}{{\rm{v}}_{1\times 1}}$ represents a 1$\times$1 convolutional layer for feature channel adjustment. ${\rm{con}}{{\rm{v}}_{3\times 3}}$ is a 3$\times$3 convolutional layer for feature fusion. $\rm{Up}$ means bilinear upsampling. Similarly, we construct a bottom-up path to flow the lowest-level feature in the encoder to the topmost feature, so that the resulting features aggregate the lower-level features. We still employ a simple feature fusion module to fuse adjacent-level features, but we use bilinear interpolation to downsample features instead of upsampling. This can be mathematically expressed as:
@@
Figure 5: Long-range Skip Connection based Aggregation architecture (our initial design).
We compare our MFAM with FPN-like architecture [16, 25, 19, 42] (Fig. 2(b)) and Long-range Skip Connection based Aggregation (LSCA) architecture illustrated in Fig. 5 to give readers a clearer understanding of the motivations of our design. FPN-like architecture utilizes the top-down path with lateral connections to obtain high-level semantic features at all stages, while its topmost and mid-scale features lack low-level information. The LSCA is the one we originally designed for fusing multi-level features. Although it could fuse multi-level features in an explicit manner, using the long skip connections between long-distance features (e.g., ${S_{4}}$ to ${S_{1}}$) for fusion will lead to a larger semantic gap and misalignment than fusion between adjacent features. Quantitative comparisons are given in Table. 2.
3.3 Recursive Alignment Module (RAM)
Inspired by [47, 48, 42], we adopt a flow-based alignment module in RAM to learn 2D transformation offsets to guide the upsampling between features with different resolutions. However, we have noticeable differences from above methods. Firstly, the aim of our RAM is to achieve spatial alignment between score maps by aligning features before the segmentation head, not for feature fusion. Secondly, the previous MFAM is complementary to our RAM, which can narrow the semantic gap between multi-scale features to prevent the offset learning from being misguided. Thirdly, we align features in a stepwise manner specifically designed for long-distance feature maps, while the above methods use their flow-based alignment module for adjacent features without large resolution differences.
@@ 3.3 Recursive Alignment Module (RAM)
In Fig. 4, we take the alignment between ${F_{4}}$ and ${F_{1}}$ as an example to compare our proposed RAM with the straightforward alignment method. As shown in Fig. 4, the straightforward method puts ${F_{4}}$ and ${F_{1}}$ into the flow-based alignment module to learn the corresponding offset, and then use the alignment function to directly align ${F_{4}}$ to ${F_{1}}$ without using intermediate features. Compared with the straightforward alignment, our designed RAM is not only more accurate but also more efficient. Here, we give computational complexity analysis. Since the complexity of other operations in RAM and the straightforward alignment method is negligible compared to the 3 $\times$ 3 convolutional operation, the overall complexity can be estimated by 3 $\times$ 3 convolutional operations. Therefore, the complexity of computing the offset between ${F_{4}}$ and ${F_{3}}$ can be expressed as $W\times H\times{C_{in}}\times{C_{out}}\times 3\times 3$, which can be simplified as $9WH{C^{2}}$. Here, $H$ and $W$ denote the height and width of ${F_{3}}$, and ${C_{in}}$ and ${C_{out}}$ represent the channels of input and output features of the convolutional layer. Since the height and width of ${F_{1}}$ are increased by 4 times compared to ${F_{3}}$, the complexity of inputting upsampled ${F_{4}}$ and ${F_{1}}$ into the 3 $\times$ 3 convolutional layer can be expressed as $4W\times 4H\times{C_{in}}\times{C_{out}}\times 3\times 3$. Ultimately, the ratio between the overall complexity of RAM and the straightforward alignment method is: $(9WH{C^{2}}+4\times 9WH{C^{2}}+16\times 9WH{C^{2}})/(3\times 16\times 9WH{C^{2}})=7/16$. In terms of accuracy, the quantitative comparison and corresponding analyses are given in Section. 4.2.3.
3.4 Adaptive Scores Fusion Module
@@
3.5 Multi-scale Joint Supervision
@@ 4.1 Training Protocol
Table 1: Comparison with the baseline method. MSL: Multi-Stage Loss. MJS: Multi-scale Joint Supervision. OHEM: Online Hard Example Mining.
@@
Table 2: Comparison with different multi-level feature aggregation methods. FPN-like: the FPN-like Network shown in Fig. 2(b). LSCA: the Long-range Skip Connection based Aggregation Architecture in Fig. 5.
@@
Table 3: Ablation study on RAM. Bilinear: bilinear interpolation. SA: Straightforward Alignment.
@@
Table 4: Comparison between three multi-scale scores fusion methods: Average, Maximum, and ASFM.
@@
Table 5: Ablation study on importance of each component. ”-”: Indicates that the corresponding experiment directly utilizes the features of the encoder for alignment and prediction. Bilinear: bilinear interpolation. Cancat: the channel concatenation operation for feature fusion.
@@
Table 6: Ablation Study for booster training strategy. OHEM: Online Hard Example Mining.
@@
4.2 Ablation Studies
4.2.1 Comparison with the Baseline Methods
Here, we choose a modified version of FPN network adapted for segmentation tasks as the baseline method (named as ’native FPN’ in Table. 1), which aggregates the multi-level outputs of FPN to the highest resolution feature for single-scale prediction [25, 19, 42]. For a fair comparison, we also adopt two atrous convolutional layers to the backbone of the baseline, following the way described in Section. 3.2. Furthermore, we add three additional segmentation heads at different stages of the baseline to compute auxiliary loss to boost network training, which is named Multi-Stage Loss (MSL). Notably, by inferring on multi-scale features, our MFARANet can calculate MSL without adding additional segmentation head. As shown in Table. 1, through the above operations, we improve the accuracy of the native FPN from 72.3% to 74.7% of the improved FPN with a slight decrease in inference speed. Using the same backbone and MSL, our method can achieve 77.0% mIoU, which is 2.3% higher than the strong baseline. At the same time, the computational complexity is reduced by 10.5 GFLOPs, achieving a faster inference speed. Furthermore, by combining MJS and OHEM to boost the training process, our method further improves the accuracy to 78.2%, which is 5.9% and 3.5% higher than native FPN and improved FPN, respectively. Qualitative segmentation comparison results between our method and the baseline on Cityscapes validation set are shown in Fig. 7. Our method shows improvement on multi-scale objects.
4.2.2 Ablation Study for MFAM
To demonstrate the superiority of MFAM, we compare it with the other two feature aggregation methods: FPN-like architecture (Fig. 2(b)) and LSCA (Fig. 5). In Table. 2, we replace the MFAM in our network with FPN and LSCA, respectively, and all other components remain unchanged. From the results, our method improves the accuracy by 0.7% and 1.6% compared to FPN and LSCA, respectively, with less than 6 GFLOPs increase in computational complexity. We attribute the superiority of our MFAM as two aspects: a) The topmost and mid-scale features of FPN lack low-level features, while each feature from MFAM aggregates full-level features of the encoder. b) We aggregate long-distance features in an information flow manner, which narrows the semantic gap compared to using long-range skip connections.
4.2.3 Ablation Study for RAM
We propose RAM to align scaled features and carry out several ablation studies to reveal its effectiveness. First, we utilize bilinear interpolation to replace RAM for upsampling, and the first row in Table. 3 has a decrease in accuracy of 1.5% compared to the third row. This shows that spatial misalignment is one of the key issues for multi-scale scores fusion and can be efficiently solved by our RAM. Second, we compare Straightforward Alignment (Fig. 4) with our proposed Recursive Alignment (Fig. 4), and our RAM improves the accuracy and inference speed by 0.5% and 2.6 FPS, respectively. We attribute this to adopting the stepwise alignment to replace direct alignment between long-distance features. Adjacent features have less spatial misalignment due to smaller scale differences compared to long-range features, leading to more accurate offsets learning. Computational complexity analysis has been given in Section. 3.3.
[…]
Table 7: Ablation for Multi-scale predictions. ”-”: The corresponding score map and weight map are discarded during training and inference phase. The proposed MJS + OHEM is used to supervise the training of each experiment.
@@ 4.2.3 Ablation Study for RAM
Table 8: Experiments for the scale selection during the inference to further reduce model complexity. All ablations are trained on the complete architecture with four scales and discarded certain scales only during the inference phase. The results reaching the best trade-off are bolded.
@@
4.2.4 Ablation Study for ASFM
We design ASFM to fuse multi-scale scores in an adaptive manner. Three fusion methods are compared in Table. 4: Average, Maximum (Max), and ASFM. When using Average or Max operations, the score maps at each scale are either equally important or sparsely selected, making them a suboptimal choice. ASFM improves the segmentation accuracy by more than 1.0% over the other two methods, demonstrating the effectiveness of this adaptive fusion strategy. Compared with the promotion of accuracy, the increase in computational complexity is acceptable, as the inference speed is only reduced by around 5 FPS and remains above the real-time speed.
4.2.5 Importance of each Module
In Table. 5, we conduct three experiments to reveal the importance of feature aggregation, spatial alignment and adaptive score-level fusion. In the first row, we directly input the features from each level of the encoder into subsequent modules. This results in a significant reduction in accuracy of 2.6%, suggesting that aggregating multi-level features at each scale can benefit spatial alignment and inference. In the second experiment, we utilize bilinear interpolation to replace RAM for upsampling low-resolution features, which decreases the accuracy by 1.5%. This result reveals that spatial misalignment is the key factor limiting scores fusion, and our RAM can effectively address this problem. Finally, we use concatenation operation followed by a 3 $\times$ 3 convolutional layer to fuse multi-scale features first, and then use a segmentation head to obtain the prediction. We conduct this experiment to compare the commonly used feature-level fusion with our in-network score-level fusion. Our ASFM reduces the computational complexity by 17.2 GFLOPs while improving the accuracy by 0.6%, which proves the effectiveness and efficiency of the adaptive score-level fusion mechanism.
4.2.6 Ablation Study for Booster Training Strategy